40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 1 08:50:48 2022, yuta, Summary, LSC, FPMI with REFL/AS55 trial 
    Reply  Thu Jul 14 00:10:08 2022, yuta, Summary, LSC, FPMI with REFL/AS55 trial continued AS55_I.pngAS55_Q.pngREFL55_I.pngREFL55_Q.png
       Reply  Fri Jul 15 19:13:22 2022, Paco, Summary, LSC, FPMI with REFL/AS55 demod phase adjust 
          Reply  Fri Jul 15 22:36:04 2022, rana, Summary, LSC, FPMI with REFL/AS55 demod phase adjust 
          Reply  Mon Jul 18 16:39:07 2022, Paco, Summary, LSC, FPMI locking procedure using REFL55 and AS55 DARM_07_18_2022_FMPI.pdfCARM_07_18_2022_FPMI.pdfMICH_07_18_2022_FPMI.pdffpmi_darm_nb_2022_07.pdf
             Reply  Mon Jul 18 21:41:42 2022, Anchal, Summary, LSC, FPMI locking procedure using REFL55 and AS55 
             Reply  Tue Aug 9 19:54:31 2022, yuta, Summary, LSC, FPMI locking tonight 
                Reply  Thu Aug 18 14:49:35 2022, Yehonathan, Summary, LSC, FPMI Sensitivity Sensitivity_Plot_1344133503.pdf
                   Reply  Thu Aug 18 18:10:49 2022, Koji, Summary, LSC, FPMI Sensitivity 
Message ID: 17002     Entry time: Thu Jul 14 00:10:08 2022     In reply to: 16968     Reply to this: 17007
Author: yuta 
Type: Summary 
Category: LSC 
Subject: FPMI with REFL/AS55 trial continued 

[Paco, Koji, Yuta]

We managed to lock MICH using REFL55_Q by setting the demodulation phases and offsets right.
The following is the current FPMI locking configuration we achieved so far.

DARM: POX11_I / gain 0.007 / 0.5*ETMX-0.5*ETMY (or 1*ETMX) / UGF of ~100 Hz
CARM: POY11_I / gain 0.018 / 1*MC2 / UGF of ~200 Hz
MICH: REFL55_Q / gain -10 / 0.5*BS / UGF of ~30 Hz

Transitioning DARM error signal from POX11_I to 0.5*POX11_I+0.5*POY11_I was possible with FM4 filter off in DARM filter bank, but not to AS55_Q yet.

REFL55 and AS55 demodulation phase tuning:
 - We found that both AS55 and REFL55 are contaminated by large non-MICH signal, by making a ASDC vs RF plot (see 40m/16929).
 - After both arms are locked with POX and POY, MICH was locked with AS55_Q. ASDC was minimized by putting an offset to MICH filter.
 - With this, REFL55 offsets were zeroed and demodulation phase was tuned to minimize REFL55_Q.
 - Locked MICH with REFL55_Q, and did the same thing for AS55_Q.
 - Resulting ASDC vs RF plots were attached. REFL55_Q now looks great, but REFL55_I and AS55 are noisy (due to signals from the arms?).

Jupyter notebook: https://git.ligo.org/40m/scripts/-/blob/main/CAL/MICH/MICHOpticalGainCalibration.ipynb

Sensing matrix:
 - With FPMI locked using POX/POY, DARM and CARM lines were injected at around 300 Hz to measure the sensing gains. For line injection, C1:CAL-SENSMAT was used, but for the demodulation we used a script. The following is the result.

 Sensors              DARM (ETMX)         CARM (MC2)        
C1:LSC-AS55_I_ERR    3.10e+00 (-34.1143 deg)    1.09e+01 (-14.907 deg)    
C1:LSC-AS55_Q_ERR    9.96e-01 (-33.9848 deg)    3.30e+00 (-27.9468 deg)    
C1:LSC-REFL55_I_ERR    6.75e+00 (-33.7723 deg)    2.92e+01 (-34.0958 deg)    
C1:LSC-REFL55_Q_ERR    7.07e-01 (-33.4296 deg)    3.08e+00 (-33.4437 deg)    
C1:LSC-POX11_I_ERR    3.97e+00 (-33.9164 deg)    1.51e+01 (-30.7586 deg)    
C1:LSC-POY11_I_ERR    6.25e-02 (-20.3946 deg)    3.59e+00 (38.4207 deg)

Jupyter notebook: https://git.ligo.org/40m/scripts/-/blob/main/CAL/SensingMatrix/MeasureSensMat.ipynb

 - By taking the ratios of POX11_I and AS55_Q for DARM, POY11_I and REFL55_I for CARM, we tried to find the correct gains for REFL55 and AS55 for DARM and CARM. x3.96 more gain for AS55_Q than POX11_I and x0.123 less gain for REFL55_I than POY11_I.

Next:
 - Try locking the arms with no triggering, and then try locking FPMI with REFL/AS without triggering. No FM4 for this, since FM4 kills gain margin.
 - Lock single arm with AS55_Q and make a noise budget. Make sure to misalign ITMX(Y) completely when locking Y(X)arm.
 - Lock single arm with REFL55_I and make a noise budget.
 - Repeat Xarm noise budget with Yarm locked with POY11_I and MC2 (40m/16975).
 - Check IMC to reduce frequency noise (40m/17001)

Attachment 1: AS55_I.png  88 kB  Uploaded Thu Jul 14 01:15:14 2022  | Hide | Hide all
AS55_I.png
Attachment 2: AS55_Q.png  77 kB  Uploaded Thu Jul 14 01:15:22 2022  | Hide | Hide all
AS55_Q.png
Attachment 3: REFL55_I.png  113 kB  Uploaded Thu Jul 14 01:15:29 2022  | Hide | Hide all
REFL55_I.png
Attachment 4: REFL55_Q.png  72 kB  Uploaded Thu Jul 14 01:15:39 2022  | Hide | Hide all
REFL55_Q.png
ELOG V3.1.3-